OpenStudy (anonymous):

Would you step me through solving this equation step by step? Solving for (f) I get +-5 on the calc but I can't replicate by hand. Help!

OpenStudy (anonymous):

what is the question +-5?

OpenStudy (anonymous):

\[1000000/\sqrt(10000^2+(1/(i*2*\pi*f*31.83E-9))^2) = 1\] Sorry I forgot the equals one.